OpenStudy (anonymous):

prove the identity: 1 over sec x + tan x = 1 - sin x over cos x

OpenStudy (anonymous):

\[\left( 1 \over secx + tanx \right) = \left( 1-sinx \over \cos x \right)\]

OpenStudy (anonymous):

LHS=cosx / (1+sinx) LHS/RHS = (cos^2 x)/(1-sin^2 x)=cos^2 x / cos^2 x =1 so LHS=RHS

OpenStudy (anonymous):

cosx / (1+sinx) ?

OpenStudy (anonymous):

yeah that is LHS

OpenStudy (anonymous):

multiply the numerator and denominator by cosx

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

I don't get LHS/RHS? what's going on hehe

OpenStudy (anonymous):

divide LHS by RHS and found it is 1. so we can say LHS=RHS

OpenStudy (anonymous):

(1-sinx) times (1+sinx) is one?

OpenStudy (anonymous):

(1-sinx) times (1+sinx) =